OpenStudy (kamille):

Hello, can someone help me? I have a reaction: 2Mg+O2->2MgO Also, I know that mass of Mg is 9g mass of O2 is 8g What is the mass of MgO?

OpenStudy (yrelhan4):

find the moles f each mg and o2. check which is the limiting reagent. that many moles of mgo will be formed. also note x moles of mg will give x moles of mgo. but x moles of o2 will give 2x oles of mgo.

OpenStudy (anonymous):

1. Find moles of Mg and O2 Mg → 9/24.3 = .37 mol Mg O2 → 8/16.0 = .50 mol O2 2. Find the limiting reactant (moles of each substance divided by coefficient) Mg → .37mol ÷ 2 mol = .19 O2 → .50 mol ÷ 1 mol = .5 Limiting reactant (smallest number) = Mg 3. Use moles of Mg to find how many moles of MgO formed \[\frac{ 2 mol Mg }{ 2 mol MgO } = \frac{ .5 mol O2 }{ x }\] x = .5 mol MgO formed 4. Convert moles of MgO to grams MgO → .5(24.3g) + .5(16.0g) = 20.15g Answer: 20.15 grams MgO